The following data apply to questions 3 through 10. Sebastian Company, which manufactures electrical switches, uses a standard cost system and carries all inventories at standard. The standard manufacturing overhead costs per switch are based on direct labor hours and are shown below: Variable overhead (5 hours @ $12 per direct manufacturing labor hour) $ 60 Fixed overhead (5 hours @ $15* per direct manufacturing labor hour) 75 Total overhead per switch $135 *Based on capacity of 200,000 direct manufacturing labor hours per month. The following information is available for the month of December:  46,000 switches were produced although 40,000 switches were scheduled to be produced.  225,000 direct manufacturing labor hours were worked at a total cost of $5,625,000.  Variable manufacturing overhead costs were $2,750,000.  Fixed manufacturing overhead costs were $3,050,000. 3. The variable overhead spending variance for December was a. $50,000 U. b. $350,000 U. c. $10,000 F. d. $60,000 F. 4. The variable manufacturing overhead efficiency variance for December was a. $50,000 U. b. $350,000 U. c. $10,000 F. d. $60,000 F. 5. The total variable manufacturing overhead variance was a. $10,000 F. b. $10,000 U. c. $110,000 U. d. $110,000 F. 6. The fixed manufacturing overhead spending variance for December was a. $450,000 F. b. $400,000 F. c. $50,000 U. d. $775,000 F. 7. The fixed overhead production-volume variance for December was a. $450,000 F. b. $400,000 F. c. $50,000 U. d. $775,000 F. 8. What amount should be credited to the Allocated Manufacturing Overhead Control account for the month of December? a. $6,210,000 b. $5,800,000 c. $5,760,000 d. $5,700,000 9. The manufacturing overhead flexible-budget variance for December was a. $10,000 F. b. $40,000 U. c. $50,000 U. d. $100,000 U. 10. The manufacturing overhead spending variance for December was a. $10,000 F. b. $40,000 U. c. $50,000 U. d. $100,000 U.
Answer #1

ANSWER

3).

Variable Overhead spending variance = Actual variable overhead - Budgeted Variable overhead
= $2750000 - (40000 x 5 x $12) = $350000 (U)

Answer is b. $350000 (U)

4).

Variable overhead efficiency variance = (Actual hours - Standard hours) x Standard rate
= (225000 - 230000) x $12 = $60000 (F)

Standard hours = 46000 x 5 = 230000 hours

Answer is d. $60000 (F)

6).

Fixed overhead spending variance = Actual fixed overhead - Budgeted fixed overhead
= $3050000 - (200000 x $15) = $50000 (U)

Answer is c. $50000 (U)

7).

Fixed overhead volume variance = (Standard hours - Capacity hours) x Overhead rate
= (230000 - 200000) x $15 = $450000 (F)

Answer is a. $450000 (F)

8).
Overhead allocated = 230000 x $12 + 230000 x $15 = $6210000

Answer is a. $6210000